OpenStudy (anonymous):

F(x+h) - F(x)/h F(x)=x^2-7

OpenStudy (alekos):

F(x+h) = (x+h)^2 - 7 = x^2 + 2hx + h^2 - 7 and we know that F(x) = x^2 - 7 so I'll let you work out F(x+h) - F(x)

OpenStudy (alekos):

then we divide the result by h

OpenStudy (anonymous):

thats the same question

OpenStudy (alekos):

what do you get?

OpenStudy (anonymous):

2x+h-14

OpenStudy (anonymous):

@alekos

OpenStudy (alekos):

not quite. It's 2x + h

OpenStudy (alekos):

you need to watch the signs

OpenStudy (anonymous):

Okay thnk you

OpenStudy (alekos):

no problem